the height of a right triangle is 6 inches longer than the base. The area of the triangle is 108 sq. in. Find the measure of the triangle's base and height?
Answer by cleomenius(959)

Area = 1/2h*b
216 = (x)(x+6)
216 = x^2 + 6x
x^2 + 6x -216 = 0
(x - 12) (x + 18_
x = 12 inches, the base.
x = 18 inches, the height.
108 = 1/2(12)(18)
This does check.
Cleomenius.